OpenStudy (anonymous):

If a spherical baseball has a diameter of 7 cm, calculate the surface area and volume of the ball. (GEOMETRY!)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

AND the volume too!!

OpenStudy (anonymous):

@Directrix @Sjano

OpenStudy (anonymous):

@Preetha

hartnn (hartnn):

diameter = 7, so whats the radius ?

hartnn (hartnn):

radius is half the diameter.

OpenStudy (anonymous):

3.5

hartnn (hartnn):

right , now surface area is \(\huge 4\pi r^2\)

OpenStudy (anonymous):

49 pie?

hartnn (hartnn):

yup, 49pi .....if u want answer without pi, u can put pi=3.14 or 22/7

OpenStudy (anonymous):

with pie and for volume as well

hartnn (hartnn):

Volume = \(\huge \frac{4}{3}\pi r^3\)

OpenStudy (anonymous):

I got 42.875

hartnn (hartnn):

i m getting 57.16 pi

hartnn (hartnn):

57.16 pi cm^3

OpenStudy (anonymous):

Oh...wait..

Directrix (directrix):

V = 4/3 * pi* (3.5)^3 SA = 4 * pi * (3.5)^2

OpenStudy (anonymous):

I typed it in wrong, I got 57.16 too

hartnn (hartnn):

then its correct 57.16 pi cm^3 to be complete....

Directrix (directrix):

V = (343/6) pi cubic centimeters SA = 49 pi square centimeters